Take the density of the crown to be rho_c. What is
Answer #1

Part A) We know that W = \rho Vg

Actaual weight of crown   W_{actual} = \rho_{c} Vg

Apprant weight of crown = Actual weight of crown - water displaced

W_{apparent} = \rho_{c} Vg- \rho _{w}Vg

\frac{W_{apparent}}{W_{actual}} = \frac{\rho_{c} Vg- \rho _{w}Vg}{\rho_{c} Vg}= 1-\frac{\rho _{w}}{\rho _{c}}

Part B

\frac{W_{apparent}}{W_{actual}} =\frac{4.5}{5}= 1-\frac{\rho _{w}}{\rho _{c}}

\frac{\rho _{w}}{\rho _{c}}= 1- \frac{4.5}{5} = 0.1

\rho _{c}= \frac{\rho _{w}}{0.1} = \frac{1/ccg}{0.1} = 10 g/cc

As \rho _{c}< \rho _{gold}   so crown is not made of pure gold.